OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -4, -16, -64, -256, ...

OpenStudy (jdoe0001):

any ideas on what the "common ratio" or multiplier is? what value you'd need to multiply the value to get the next value that is

OpenStudy (anonymous):

hint: geometric

OpenStudy (jdoe0001):

\(\large \bf a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1}\qquad \begin{array}{llll} a_1=\textit{1st term of sequence}\\ r=\textit{common ratio or multiplier} \end{array}\)

OpenStudy (anonymous):

Wait, so which of the options would it be? a. an = 4 • -4n b. an = 4 • -4n + 1 c. an = -4 • 4n d. an = -4 • 4n - 1

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well.. what did you get for the multiplier?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

so... -4 * -4 = -16 ? because from -4 to -16.... if we use -4.... it should be the next term, which is -16

OpenStudy (anonymous):

Whoops. So it's +4 so the answer is d. an = -4 • 4n - 1 right? @jdoe0001

OpenStudy (jdoe0001):

r = +4. yeap... thus \(\large \bf a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1}\implies a_{\color{red}{ n}}=-4\cdot 4^{{\color{red}{ n}}-1}\)

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

yw